A cat ran 30 miles per hour. AT this rate, how many minutes would it take the cat to run 1.5 miles?
A-0.05
B-3
C-20
D-45

Answer by stanbon(75887)

A cat ran 30 miles per hour. AT this rate, how many minutes would it take the cat to run 1.5 miles?
----
Use a proportion::
time/dist = time/distance
---
t/1.5 = 1/30
----
t = 1.5/30
t = 15/300
t = 1/20 = 5/100 = 0.05
------------
Cheers,
stan H.
--------------
A-0.05
B-3
C-20
D-45